The suetterl package^{∗} for use with L^{A}TEX2e

Gerd Neugebauer Mainzer Str. 16 56321 Rhens

             gerd@informatik.uni-koblenz.de

1996/07/18

1 Introduction

The font suetterl provides a suetterlin kind of font. This fonts are the handwritten variants of the gothic fonts. The sutterl and the schwell font have been written by B. Ludewig. They can be found on the CTAN in the directory tex-archive/fonts/gothic/sueterl. This package provides means to use this font.

2 Usage

This file can be used as a package by placing its name in the argument of \usepackage. Afterwards the font family suetterl is defined. This could also have been done by providing a font definition file.

The font definitions in this file scale down the original fonts to let L^{A}TEX choose the right baselineskip.

\suetterlin The command \suetterlin changes the current font family to suetterl and the encoding to T1. Usually this should be used in a TEX group only since the macros \s is redefined as well.

\textsuetterlin The command \textsuetterlin typesets its argument in the suetterl font.

3 Implementation

First we declare a new font family for the suetterl font.

 1 \DeclareFontFamily{T1}{suetterl}{}

This font is only available in the normal shape. Here we can get the desired font by (silently) scaling the only present suett15. The appropriate factor seems to be magstep 2 = 1.44.

 2 \DeclareFontShape{T1}{suetterl}{m}{n}{<->s*[0.8]suet14}{}
 3 \DeclareFontShape{T1}{suetterl}{m}{it}{<->s*[0.7]schwell}{}

∗This file has version number v1.2, dated 1996/07/18.

Now we define font changing command.

\sutterlin The macro \suetterlin selects the suetterl family. The macro \s is used for the short s.

4 \DeclareRobustCommand\suetterlin{%
5 \renewcommand\s{\symbol{28}}%
6 \fontfamily{suetterl}%
7 \fontencoding{T1}%
8 \selectfont}

\s Usually \s produces a normal s.

9 \newcommand\s{s}

\textsutterlin The macro \textsuetterlin typesets is argument in the sutterl font.

10 \newcommand\textsuetterlin[1]{\begingroup\suetterlin #1\endgroup}